# Face Forgery Detection with Elaborate Backbone

### v1

@@ 3.2 FFD with Pre-trained Backbones
We abstract a traditional pipeline of existing FFD methods to highlight the vital role of backbones, as shown in Figure 4(top). These methods directly use backbones pre-trained with SL on ImageNet, implement fine-tuning networks or strategies to equip the backbone for better generalization, and apply an empirical classification threshold ($0.5$) to output the final discriminant result (real or fake).
@@
In our comprehensive comparison (refer to Section 4.2), we statistically find that backbones pre-trained with SSL outperform those with SL under the same network architecture, and Transformer-based backbones outperform CNN-based ones with the same learning approach. These results are consistent with recent research trends observed across various tasks. Notably, BEiT [25, 26] achieves superior performance, benefiting from its ability to independently represent visual tokens of image patches, thus better extracting differential cues between facial local components.
@@
Since most mainstream backbones are pre-trained on ImageNet, which contains content significantly different from facial images, this training data gap will also restrict the performance of general-purpose backbones in facial-related tasks, particularly FFD, which requires identifying subtle forgery cues within complex facial features. Furthermore, compared to forgery faces, we can easily collect large-scale and diverse real-face datasets [103, 104]. Although these datasets may lack labels, we can still leverage SSL with appropriate pretext tasks to pre-train a robust face representation backbone with face-specific knowledge tailored to facial tasks. In the technical implementation of the backbone, we adopt ViT as the network architecture and employ SSL methods such as MoCo v3 [90], MAE [24], and BEiT v2 [25], which are concise and classical in CL and MIM.
@@
As expected, the backbone pre-trained on real-face datasets yields superior FFD generalization compared to one on ImageNet, as evident in SSL-Face and SSL-1k&Face of Table I. Notably, MoCo v3 and BEiT v2 are superior to MAE, whose performance is relatively weaker. We attribute these results to the close relationship between FFD core challenges and the backbone’s pretext tasks. On the one hand, forgery clues typically manifest in the local components of fake faces (e.g., eyebrows, eyes, lips) and the contrast between facial components, as depicted in Figure 2. On the other hand, MoCo v3, with its CL pretext task that compares the similarity of high-level features from differently augmented images, can better focus on forgery cues in local regions between these images. Meanwhile, with a classification pretext task based on encoding image patches, BEiT v2 excels at discerning the authenticity of local facial components. The excellent performance of MoCo v3 and BEiT v2 further demonstrates the indispensability of equipping backbones with both general and specific knowledge through appropriate pre-training methods to enhance FFD generalization.
@@
Based on our comprehensive study of backbones in the pre-training stage, we revitalize the traditional FFD pipeline by integrating SSL on real-face datasets, a competitive backbone framework, and an uncertainty-based threshold optimization mechanism in the pre-training, fine-tuning, and inference stages. As shown in Figure 4(bottom), our revitalized FFD pipeline enables a more proficient and promising step-by-step enhancement of the backbone’s baseline and task-specific capabilities, as well as greater reliability in FFD discriminant results. Additionally, given the superior performance of BEiT v2, we use its SSL method and ViT architecture as the representative pre-training configuration in our pipeline, as shown in Figure 4(left-bottom). Next, we introduce our competitive backbone fine-tuning framework.
@@ 3.3 Competitive Backbone Fine-tuning Framework
Although the backbone pre-trained on real-face datasets better represents facial features, employing promising fine-tuning methods to equip the backbone with expert knowledge from forgery faces remains crucial for FFD. Hence, we further analyze the inherent characteristics of FFD problem. Humans often identify forged faces by noticing semantic repetition or inconsistencies in facial components and contrast in appearance between them, as illustrated in Figure 2. Correspondingly, the apparent and implicit forgery cues in the latent space of forged faces are even more diverse and entangled, making it nearly impossible to explicitly disentangle these two types of cues due to the lack of available definitive constraints. Therefore, we consider leveraging two backbones to competitively mine different forgery cues in the latent space to enhance generalization.
Inspired by MoCo v3 with a dual-branch Siamese architecture, where the pre-trained backbone effectively extracts facial component features and achieves excellent FFD performance, we propose building a dual-branch framework that cultivates two backbones to extract different forgery cues within a competitive adversarial learning mechanism. However, MoCo v3 primarily constrains the backbones to represent the semantic similarity of two differently augmented images, which involve changes in both high-level (e.g., crop, resize) and low-level (e.g., color distortion, Gaussian blur) features. These different augmentations of the same image may potentially lead to the loss or introduction of subtle forgery traces related to low-level features, such as artifacts and appearance. Thus, we remove the image augmentations and contrastive loss while retaining only the dual-branch architecture and incorporating cross-entropy loss $\mathcal{L}_{ce}$ for FFD. Furthermore, we fuse the output features of both backbones to enable their interaction, thereby forming our basic dual-branch backbone fine-tuning architecture.
Since the dual branches have the same inputs (without different augmented images) and loss functions ($\mathcal{L}_{ce}$), the features extracted by both backbones will inevitably be similar. Thus, this basic dual-branch architecture alone cannot make both backbones competitively learn different forgery cues. To address this, we introduce a feature metric as a decorrelation constraint to minimize the correlation between features extracted by both backbones, enabling competitive learning within the dual-branch architecture.
@@
At this point, the dual-branch architecture with decorrelation constraint is preliminarily established, as shown in Figure 4(middle-bottom), namely competitive backbone fine-tuning framework. It includes two unshared backbones processing the same face image, feature fusion module, and three $\mathcal{L}_{ce}$, which constrain both backbones to competitively mine different forgery cues to enhance the main branch’s generalization. $\mathcal{L}_{dec}$ and $\mathcal{L}_{ce}$ are defined as follows:
$\vspace{-10pt}\mathcal{L}_{dec}$ $=\frac{\sum_{i=1}^{C}(f_{M}^{i}-\overline{f}_{M})(f_{A}^{i}-\overline{f}_{A})}{\sqrt{\sum_{i=1}^{C}(f_{M}^{i}-\overline{f}_{M})^{2}\sum_{i=1}^{C}(f_{A}^{i}-\overline{f}_{A})^{2}}},$ (1)
$\mathcal{L}_{ce}$ $=-\left(y\log(p)+(1-y)\log(1-p)\right),\vspace{-10pt}$ (2)
where $f$ represents the feature vectors from both main ($M$) and auxiliary ($A$) backbones with dimensions $C$, while $y$ and $p$ are the label and predicted probability, respectively.
@@
Standard classification networks typically use Softmax in their final layer to convert the continuous output values into class probabilities. However, Softmax tends to exaggerate the probabilities of predicted classes due to its exponential function, which cannot accurately reflect the model’s prediction confidence. Despite minimizing the $\mathcal{L}_{ce}$ of predicted classes during training, the classification model lacks awareness of its prediction confidence. To overcome this issue, uncertainty estimation methods have been proposed to assess the model’s confidence in its predictions [37]. Particularly, Evidence Deep Learning (EDL) simultaneously predicts a model’s class probability and its uncertainty by assuming that the predicted class probabilities follow a prior Dirichlet distribution [105, 106, 107]. Therefore, we utilize the uncertainties ($u$) from both branches as weights ($w$) to fuse their feature vectors, as shown in Figure 5, enabling flexible and dynamic competition between both backbones. The fusion process is as follows:
@@
$-\left(1-\lambda_{t}\right)\sum_{i\in\left\{\hat{y}\neq y\right\}}\left(1-p\right)\log\left(u\right),\vspace{-10pt}$ (8)
@@
As shown in Figure 4(middle-bottom), we replace the $\mathcal{L}_{ce}$ with $\mathcal{L}_{EDL}$ and $\mathcal{L}_{EUC}$ in both the main and auxiliary branches. Thus, the loss function $\mathcal{L}$ of our competitive backbone fine-tuning framework for FFD task is as follows:
$\vspace{-10pt}\mathcal{L}=$ $\mathcal{L}_{EDU}^{M}+\mathcal{L}_{EDU}^{A}+\mathcal{L}_{ce}^{F}+\mathcal{L}_{dec},\vspace{-10pt}$ (9)
@@ 3.4 FFD Inference with Optimized Threshold
In academic research, FFD models’ performance is typically evaluated using the Area Under the Receiver Operating Characteristic Curve (AUC), which measures overall results across a dataset with thresholds sampled at fixed intervals without needing explicit real or fake outputs [108]. However, in practical FFD applications, a threshold $\tau$ is required as an anchor value to classify the FFD model’s continuous probability outputs ($p\in[0,1]$) into discrete results, such as $p\geq\tau$ indicates fake and $p<\tau$ indicates real. This shows that the value of $\tau$ is crucial for the classification results, and it also determines the model’s performance on evaluation metrics such as ACC, Precision, Recall, and F1-score [109]. However, most FFD works empirically set $\tau=0.5$, it is usually not optimal, as each model’s optimal threshold is different and depends on various factors such as its architecture, training approach, and dataset.
Due to the overfitting performance of FFD models within the training dataset, their predicted probabilities $p$ are typically polarized, with real being $p\approx 0$ and fake being $p\approx 1$. In such cases, $\tau$ has little impact on within-dataset accuracy. However, given the diverse forgery cues across different cross-datasets, where $p$ values are more evenly distributed, this leads to $\tau$ becoming highly sensitive for FFD model prediction accuracy. Therefore, we strive for a threshold selection mechanism by selecting the optimal threshold $\tau_{ot}$ on one cross-dataset and validating its efficacy on others.
Specifically, we iterate through the predicted probabilities of all faces in one cross-dataset, treating each probability as a potential threshold and calculating the corresponding ACC for the dataset. The threshold with the highest ACC is determined to be $\tau_{ot}$. The calculation of $\tau_{ot}$ is as follows:
[…]
$\text{ACC}(\tau)$ $=\frac{1}{N}\sum_{i=1}^{N}\left[\mathbb{I}(p_{i}\geq\tau)y_{i}+\mathbb{I}(p_{i}<\tau)(1-y_{i})\right],\vspace{-10pt}$ (11)
@@
where $N$ is the total number of face images, $y_{i}$ is the label for the $i$-th image (with $1$ for fake and 00 for real), and $\mathbb{I}$ is the indicator function returning $1$ if true and 00 otherwise.
Note that the method above for finding the $\tau_{ot}$ resembles a white-box scenario. When the predicted probabilities of all images are unknown but the overall ACC is available, we can still approximate the optimal threshold by searching over $\tau\in[0,1]$ with a defined step size. Moreover, the comparison of $\tau=0.5$ and $\tau=\tau_{ot}$ in Table VIII indicates that $\tau_{ot}$ from one cross-dataset can significantly improve the ACC of other datasets, underscoring the critical role of optimized thresholds in FFD applications.
As mentioned in the Uncertainty-based Fusion Module of Section 3.3, the class probabilities amplified by Softmax can cause the model to be overly confident in its predictions, even when it is wrong. To mitigate this, we propose introducing the uncertainty $u$ to adjust the probability $p$, which is crucial for improving the reliability of the predictions. Moreover, we devise a threshold optimization mechanism where dividing $p$ by $u$ yields the adjusted probability $p^{\prime}$, i.e., $p^{\prime}=p/u$, which enhances $p$ with high confidence (low uncertainty) and reduces $p$ with low confidence (high uncertainty), as shown in Figure 4(right-bottom). Based on the adjusted probability $p^{\prime}$, we recompute the optimal threshold using Equations 10 and 11 to improve the reliability of FFD models in final discriminant results. For more comparisons, please refer to Section 4.5.
@@ 4.1.1 Datasets
We conduct experiments by pre-training on real-face dataset, fine-tuning on c23 of Faceforensics++ [52] (FF++), and evaluating on multiple cross-datasets during inference. Specifically, our real-face dataset contains $550$,$000$ images from CelebA [103], CelebV-Text [110], and FFHQ [104]. FF++ contains $1$,$000$ real videos and $4$,$000$ manipulated videos obtained by four forging methods: DeepFakes [111], FaceSwap [112], Face2Face [113], and NeuralTexture [114]. Cross-datasets consist of Celeb-DF [53] (CDF), DFDC [54], and FFIW [115]. CDF includes real videos of $59$ celebrities from YouTube and fakes generated via an improved synthesis process [53]. DFDC is a large-scale face-changing video dataset, created using various Deepfake and GAN-based methods, with a total of $5$,$000$ test videos. FFIW is a challenging multi-face forgery dataset with $10$,$000$ videos.
To comprehensively evaluate the generalization of FFD models, we construct $9$ new cross-datasets generated by various mainstream GAN-based methods (e.g., DCGAN [116], StyleGAN3 [13], StarGAN2 [15], AttnGAN [117], E4S [14], and Wav2Lip [16]) and Diffusion Model-based methods (e.g., SD-1.5 [118], DiffFace [119], and Diffused Heads [120]), with the forged faces covering four types of forgeries: entire face synthesis, face attribute editing, face swapping, and face reenactment. Each cross-dataset consists of $5$,$000$ real faces and $5$,$000$ forged faces from each method.

### v2

@@ 3.2 FFD with Pre-trained Backbones
We abstract a traditional pipeline of existing FFD methods to highlight the vital role of backbones, as shown in Fig. 3(top). These methods directly use backbones pre-trained with SL on ImageNet, implement fine-tuning networks or strategies to equip the backbone for better generalization, and apply an empirical classification threshold ($0.5$) to output the final discriminant result (real or fake) [95, 76, 105].
@@
In our comprehensive comparison (refer to Section 4.2), we statistically find that backbones pre-trained with SSL outperform those with SL under the same network architecture, and Transformer-based backbones outperform CNN-based ones with the same learning approach. These results highlight the crucial role of SSL in pre-training ViT-based backbones to enhance FFD model generalization, with the representation capabilities of SSL approaches directly influencing FFD performance. Notably, BEiT [83, 109] achieves superior performance, benefiting from its ability to independently represent visual tokens of image patches, thus better extracting differential cues between facial local components.
@@
Since most mainstream backbones are pre-trained on ImageNet or other natural image datasets such as LAION-2B [92] and WIT [85], which contains content significantly different from facial images, this training data gap will also restrict the performance of general-purpose backbones in facial-related tasks, particularly FFD, which requires identifying subtle forgery cues within complex facial features. Furthermore, compared to forgery faces, we can easily collect large-scale and diverse real-face datasets [72, 53]. Although these datasets may lack labels, we can still leverage SSL with appropriate pretext tasks to pre-train a robust face representation backbone with face-specific knowledge tailored to facial tasks. In the technical implementation of the backbone, we adopt ViT as the network architecture and employ SSL methods such as MoCo v3 [17], MAE [44], and BEiT v2 [83], which are concise and classical in CL and MIM.
@@
As evident in SSL-Face and SSL-1k&Face of Table I, the backbone pre-trained on real-face datasets yields superior FFD generalization compared to one on ImageNet. Notably, MoCo v3 and BEiT v2 are superior to MAE, whose performance is relatively weaker. These results stem from the close relationship between FFD core challenges and the backbone’s pretext tasks. On the one hand, forgery cues typically manifest in the local components of fake faces (e.g., eyebrows, eyes, lips) and the contrast between facial components, involving both low-level features and semantic errors, as depicted in Fig. 4. On the other hand, MoCo v3, with its CL pretext task that compares the similarity of high-level features from differently augmented images, can better focus on forgery cues in local regions between these images. Meanwhile, with a classification pretext task based on encoding image patches, BEiT v2 excels at discerning the authenticity of local facial components. The excellent performance of MoCo v3 and BEiT v2 further demonstrates the indispensability of equipping backbones with both general and specific knowledge through appropriate SSL approaches to enhance FFD generalization.
@@
Based on our comprehensive study of backbones in the pre-training stage, we revitalize the traditional FFD pipeline by integrating SSL on real-face datasets, a competitive backbone framework, and an uncertainty-based threshold optimization mechanism in the pre-training, fine-tuning, and inference stages. As shown in Fig. 3 (bottom), our revitalized FFD pipeline enables a more proficient and promising step-by-step enhancement of the backbone’s baseline and task-specific capabilities, as well as greater reliability in FFD discriminant results. Additionally, given the superior performance of BEiT v2, we use its SSL method and ViT architecture as the representative pre-training configuration in our pipeline, as shown in Fig. 3(left-bottom).
@@ 3.3 Competitive Backbone Fine-tuning Framework
Building on the pre-trained backbone with enhanced facial representations, we further explore effective fine-tuning methods with improved generalization to identify implicit forgery cues in the represented facial features. Nevertheless, apparent and implicit forgery cues in the latent feature space of training sets are even more diverse and entangled, making it nearly impossible to explicitly disentangle these two types of cues due to the lack of available definitive constraints. In practice, humans tend to detect forgeries by noticing semantic repetition or inconsistencies in facial components and contrast in appearance between them, as illustrated in Fig. 4. Motivated by this, we consider leveraging two unshared backbones to competitively mine different forgery cues in the latent space to enhance generalization.
Referring to the commonly used dual-branch Siamese architecture in contrastive learning (e.g., MoCo), we propose building a dual-branch framework that cultivates two backbones to extract different forgery cues within a competitive adversarial learning mechanism. However, contrastive learning approaches primarily constrains the backbones to represent the semantic similarity of two differently augmented images, which involve changes in both high-level (e.g., crop, resize) and low-level (e.g., color distortion, Gaussian blur) features. These different augmentations of the same image may potentially lead to the loss or introduction of subtle forgery traces related to low-level features, such as artifacts and appearance. Thus, we remove the image augmentations and contrastive loss while retaining only the dual-branch architecture and incorporating cross-entropy loss $\mathcal{L}_{ce}$ for FFD. Furthermore, we fuse the output features of both backbones to enable their interaction, thereby forming our basic dual-branch backbone fine-tuning architecture.
Since the dual branches receive the same inputs (without distinct augmentations) and are optimized with the same loss ($\mathcal{L}_{ce}$), the two unshared backbones tend to learn similar features despite independent weight updates. To address this, we design a decorrelation constraint and an uncertainty-based fusion module to enforce dissimilarity between features extracted by the two unshared backbones, thereby enabling competitive learning within the dual-branch architecture, distinct from the joint or contrastive paradigms used in prior work [16, 45, 85].
@@
At this point, the dual-branch architecture with decorrelation constraint is preliminarily established, as shown in Fig. 3(middle-bottom), namely competitive backbone fine-tuning framework. It includes two unshared backbones processing the same face image, feature fusion module, and three $\mathcal{L}_{ce}$, which constrain both backbones to competitively mine different forgery cues to enhance the main branch’s generalization. $\mathcal{L}_{dec}$ and $\mathcal{L}_{ce}$ are defined as follows:
$\vskip-10.0pt\mathcal{L}_{dec}$ $=\frac{\sum_{i=1}^{C}(f_{M}^{i}-\overline{f}_{M})(f_{A}^{i}-\overline{f}_{A})}{\sqrt{\sum_{i=1}^{C}(f_{M}^{i}-\overline{f}_{M})^{2}\sum_{i=1}^{C}(f_{A}^{i}-\overline{f}_{A})^{2}}},$ (1)
$\mathcal{L}_{ce}$ $=-\left(y\log(p)+(1-y)\log(1-p)\right),\vskip-10.0pt$ (2)
where $f$ represents the feature vectors from both main ($M$) and auxiliary ($A$) backbones with dimensions $C$, while $y$ and $p$ are the label and predicted probability, respectively. Notably, we adopt $\texttt{CLS}]$ token as $f$, and utilizing all tokens from final Transformer layer yields similar performance, provided consistency with classification head input is maintained. Further analysis is available in Supp. File.
@@
FFD models typically use Softmax in their final layer to convert the continuous output values into class probabilities. However, Softmax tends to exaggerate the probabilities due to its exponential function, which does not accurately reflect the model’s prediction confidence. Despite minimizing the $\mathcal{L}_{ce}$ of predicted classes during training, the classification model lacks awareness of its prediction confidence. To overcome this issue, uncertainty estimation methods have been proposed to assess the model’s confidence in its predictions [55]. Particularly, Evidence Deep Learning (EDL) simultaneously predicts a model’s class probability and its uncertainty by assuming that the predicted class probabilities follow a prior Dirichlet distribution [93, 2]. Therefore, we utilize the uncertainties ($u$) from both branches as weights ($w$) to fuse their feature vectors, as shown in Fig. 5, enabling flexible and dynamic competition between both backbones. The fusion process is as follows:
@@
$-\left(1-\lambda_{t}\right)\sum_{i\in\left\{\hat{y}\neq y\right\}}\left(1-p\right)\log\left(u\right),$ (8)
@@
As shown in Fig. 3(middle-bottom), we replace the $\mathcal{L}_{ce}$ with $\mathcal{L}_{EDL}$ and $\mathcal{L}_{EUC}$ in both the main and auxiliary branches. Thus, the loss function $\mathcal{L}$ of our competitive backbone fine-tuning framework for FFD task is as follows:
$\mathcal{L}=$ $\mathcal{L}_{EDU}^{M}+\mathcal{L}_{EDU}^{A}+\mathcal{L}_{ce}^{F}+\mathcal{L}_{dec},$ (9)
@@ 3.4 FFD Inference with Optimized Threshold
In FFD applications, a threshold $\tau$ is required as an anchor value to classify the FFD model’s continuous probability outputs ($p\in[0,1]$) into discrete results, such as $p\geq\tau$ indicates fake and $p<\tau$ indicates real. Due to factors such as network architecture, training strategies, and dataset, prediction probability distributions for real and fake faces vary across FFD models. As a result, empirically setting $\tau=0.5$ is not suitable as a universal threshold for all FFD models. Particularly, FFD models exhibit limited generalization on unseen datasets, yielding predicted probabilities $p$ that are more evenly distributed between real and fake faces, so different choices of $\tau$ can greatly affect discriminative accuracy. In real-world scenarios, forged face images are produced by ever-emerging unseen manipulation algorithms, for which ground-truth labels are unknown to the detector. Therefore, estimating a relatively optimal threshold $\tau_{ot}$ to improve discriminative accuracy is both necessary and challenging.
Before estimating $\tau_{ot}$, as mentioned in the Uncertainty-based Fusion Module of Section 3.3, the prediction probability $p$ produced by the Softmax layer in most FFD models can cause the model to be overly confident in its predictions, even when it is wrong. To mitigate this, we introduce uncertainty $u$ to adjust $p$ via division, yielding $p^{\prime}=p/u$, where both $p$ and $u$ are predicted by our FFD model. Such adjustment increases $p^{\prime}$ under high confidence (low uncertainty) and decreases $p^{\prime}$ under low confidence (high uncertainty). Meanwhile, we apply $\tanh$ to map scores into $[0,1)$ via $p^{\prime}=\tanh(\lambda p/2(u+\epsilon))$, using $\epsilon=10^{-7}$ and scaling factor $\lambda=0.02$, which is further explained in Supp. File.
Accordingly, we propose an Uncertainty-based Threshold Optimization Mechanism (UTOM) that computes the optimal threshold $\tau_{ot}^{UTOM}$ on one cross-dataset and applies it as the relatively optimal classification threshold for other unseen datasets, which also applies in real-world scenarios where ground-truth labels for face images are unavailable. Specifically, we iterate through the adjusted probabilities $p^{\prime}$ of all faces in one cross-dataset, treating each probability as a potential threshold and calculating the corresponding ACC for the dataset. The threshold with the highest ACC is determined as $\tau_{ot}^{UTOM}$, and its calculation is as follows:
$\tau_{ot}^{UTOM}$ $=\underset{\tau}{\arg\max}\,\text{ACC}(\tau),$ (10)
$\text{ACC}(\tau)$ $=\frac{1}{N}\sum_{i=1}^{N}\left[\mathbb{I}(p^{\prime}_{i}\geq\tau)y_{i}+\mathbb{I}(p^{\prime}_{i}<\tau)(1-y_{i})\right],$ (11)
@@
where $N$ is the total number of face images, $y_{i}$ is the label for the $i$-th image (with $1$ for fake and 0 for real), and $\mathbb{I}$ is the indicator function returning $1$ if true and 0 otherwise. Moreover, comparison of $\tau=0.5$ and $\tau=\tau_{ot}^{UTOM}$ in Table IX indicates that $\tau_{ot}^{UTOM}$ from one cross-dataset can significantly improve ACC of other unseen datasets, underscoring the critical role of optimized thresholds in FFD applications.
@@ 4.1.1 Datasets
We conduct experiments by pre-training on real-face dataset, fine-tuning on c23 of Faceforensics++ [91] (FF++), and evaluating on cross-datasets with various forgery algorithms during inference. Specifically, our real-face dataset contains $550k$ images from CelebA [72], CelebV-Text [131], and FFHQ [53]. FF++ contains $1k$ real videos and $4k$ manipulated videos obtained by four forging methods: DeepFakes [25], FaceSwap [57], Face2Face [101], and NeuralTexture [100]. Cross-datasets consist of Celeb-DF [63] (CDF), DFDC [29], FFIW [139], DFDCP [28], and DFD [36]. CDF includes real videos of $59$ celebrities from YouTube and fakes generated via an improved synthesis process [63]. DFDC is a large-scale face-changing video dataset, created using various Deepfake and GAN-based methods, with a total of $5k$ test videos. FFIW is a challenging multi-face forgery dataset with $10k$ videos. DFDCP is a preview version of DFDC with $5k$ videos. DFD contains over $3k$ manipulated videos featuring $28$ actors across various scenes.
Furthermore, we evaluate the generalization ability of our method on DF40 [123], a new comprehensive FFD benchmark. Additionally, we construct $9$ new cross-datasets generated by $6$ mainstream GAN-based methods and $3$ Diffusion Model-based methods, encompassing four types of forgeries: entire face synthesis, face attribute editing, face swapping, and face reenactment. Each cross-dataset contains $5k$ real faces and $5k$ forged faces.